# with Mixture of Frequency Adaptation

### v3

@@ Abstract
Parameter-Efficient Fine-Tuning (PEFT) is a technique that allows us to adapt powerful Foundation Models (FMs) to diverse downstream tasks while preserving and unleashing their inherent capabilities. However, we have observed that existing PEFT methods, which are often designed with natural imagery in mind, struggle when applied to Remote Sensing (RS) scenarios. This is primarily due to their inability to handle artifact influences, a problem particularly severe in RS image features. To tackle this challenge, we introduce Earth-Adapter, the first PEFT method specifically designed for RS artifacts conquering. Earth-Adapter introduces a novel Mixture of Frequency Adaptation process that combines a Mixture of Adapter (MoA) with Discrete Fourier Transformation (DFT). By utilizing DFT, Earth-Adapter can decompose features into different frequency components, precisely separating artifacts from original features. The MoA then dynamically assigns weights to each adapter expert, allowing for the combination of features across various frequency domains. These simple-yet-effective approaches enable Earth-Adapter to more efficiently overcome the disturbances caused by artifacts than previous PEFT methods, significantly enhancing the FMs’ performance on RS scenarios. Experiments on Domain Adaptation (DA), and Domain Generalization (DG) semantic segmentation benchmarks showcase the Earth-Adapter’s effectiveness. Compared with baseline Rein, Earth-Adapter significantly improves 9.0% mIoU in DA and 3.1% mIoU in DG benchmarks. Our code will be released at Earth-Adapter.
1 Introduction
Tremendous last year breakthroughs in Foundation Models (FMs) [14, 45, 54, 40, 25, 55, 22, 3] have ignited transformative progress in Computer Vision (CV), demonstrating unprecedented capabilities across both high-level and low-level vision tasks. Based on these FMs, considerable Remote Sensing (RS) research groups have also constructed large-scale, state-of-the-art (SOTA), and impressive Geospatial Foundation Models (GFMs) [5, 51, 21, 65, 70, 43, 36, 23, 50, 38], significantly promoting the development of RS imagery interpretation field. However, when adapting these FMs to downstream tasks, how to efficiently keep and unleash the power of FMs becomes a key problem. Under this context, Parameter-Efficient Fine-Tuning (PEFT) methods [31, 37, 28, 8, 20] become the pivotal method due to their superior parameter-performance trade-off.
@@ 1 Introduction
Existing PEFT methods designed for language or nature imagery have made great progress with excellent works, such as LoRA [31], Visual Prompt Tuning (VPT) [37]. However, most of them show performance drops in RS cross-domain experiments. We argue the main reason lies in the influences of artifacts in the features. As shown in Figure 1 (a), we show the PCA visualization of features of DINOv2-L, which contains obvious redundant artifacts. Although a part of existing PEFT methods also are shown to be capable of handling artifacts [12], they still underestimate the artifact disturbances in RS imagery. The difference in the distribution of artifacts between natural and RS images is significant. In natural images, artifacts typically surround foreground objects [12], such as humans or animals, and the disturbances they cause are relatively limited. In contrast, RS images, due to their overhead perspective, lack centralized subjects and contain multiple coexisting multi-scale targets. For example, a single RS image may simultaneously include large-scale agricultural regions and fragmented road networks. As a result, artifacts are almost situated everywhere in RS images shown in Figure 1 (a), causing severe interference to domain-invariant representation learning, which is pivotal for cross-domain scenarios.
@@
In our view, to address the influence of artifacts, one of the core ideas is to ‘divide-and-conquer’. Drawing on the principle that high-frequency (HF) signals encode local details (e.g., edge, contour) while low-frequency (LF) signals encode global structure (e.g., object surface), we leverage Discrete Fourier Transformation (DFT) as the ‘divide’ step. As illustrated in Figure 1 (b), we successfully leverage the DFT effectively to isolate artifacts into HF signals and splits coherent global features into LF signals. Subsequently, to ‘Conquer’ artifact-related issues, we adopt a Mixture of Adapter (MoA) architecture for PEFT design. The MoA router dynamically assigns weights to adapter experts, selecting the most relevant features for processing in each iteration. In this way, all features are aggregated into optimized representations, mitigating artifacts disturbances.
Based on our motivation and findings, we further present Earth-Adapter, a simple-yet-effective RS-specialized PEFT method designed for RS artifact conquering. Specifically, Earth-Adapter consists of a novel Mixture of Frequency Adapter architecture consisting of Mixture of Adapter (MoA) with Discrete Fourier Transformation (DFT). Earth-Adapter first disentangles geospatial features by DFT, and leverages multiple experts to conquer artifacts by adaptively aggregating and learning features with suitable frequency. In this fashion, Earth-Adapter can effectively alleviate the artifacts’ disturbances for robust domain-invariant feature learning. Then gradually introduces FMs to cross the geospatial gaps from the frequency perspective.
@@
In experiments, we meticulously test Earth-Adapter on Domain Adaptation (DA), and Domain Generalization (DG) semantic segmentation tasks across five widely-used datasets. Comparing with existing PEFT methods, Earth-Adapter achieves State-Of-The-Art (SOTA) performance on DA and DG benchmarks. In DA and DG benchmarks, Earth-Adapter outperforms our baseline Rein with 8.9% and 3.2% mIoU, respectively. Additionally, we also conduct in-depth analysis of the potential of Earth-Adapter. The core contributions can be summarized as follows:
  • We propose a novel perspective to alleviate artifact-disturbance problems in RS imagery by frequency-guided divide-and-conquer strategy.
  • We introduce a simple-yet-effective PEFT method, Earth-Adapter. It is also the first RS-specialized PEFT approach designed for cross-domain semantic segmentation tasks.
  • Extensive comparisons validate the SOTA performance of Earth-Adapter on both DA and DG benchmarks. More ablation studies also demonstrate the high efficiency, robustness, and explainability of Earth-Adapter.
2 Related Work
2.1 RS Cross-Domain Semantic Segmentation
@@
As a classical and fundamental task in the realm of computer vision, semantic segmentation [46, 57] not merely holds profound practical significance, but also provides an interface for researchers to dissect the high-level semantic representation of images. With the growing need for applying models to out-of-domain data, cross-domain segmentation, including domain adaptation (DA) [29, 30, 19, 41, 33] and domain generalization (DG) [67, 21], has gained prominence. DA and DG are the two main paradigms, where DA aims to adapt a model from a labeled source domain to an unlabeled target domain, while DG focuses on training a model on only the source domain to generalize to unseen target domains. In the RS field, cross-domain task is crucial due to the diversity of data, and lots of DA methods have been applied to improve the generalization ability of segmentation models, with applications in urban planning [56], land resource management [60], and environmental protection [58]. Compared with DA, the number of DG research [34, 42] in RS is much less than DA. Although these works [72, 35, 16] make contributions, most of them only focus on specialized models rather than leveraging the power of FMs, leading to limited cross-domain capabilities.
2.2 Geospatial Foundation Models
The emergence of Large Language Models (LLMs) [4] has prompted a shift in computer vision, leading to the development of Multimodal Large Language Models (MLLMs) [48, 54, 44, 9] and Vision Foundation Models (VFMs) [45, 52, 24, 40]. FMs [3] in RS called GFMs [64, 26, 63, 62, 6] also can be divided into MLLMs and VFMs. Among MLLMs, RemoteCLIP [43] is the first exploration, supporting classification, retrieval, and object counting through a contrastive learning framework based on CLIP [54]. MetaEarth [70] advances diffusion models to generate multi-resolution RS images for any specified region based on both image and text conditions. Skysense [23] captures spatiotemporal information and performs well across diverse scenarios. Although these FMs obtain strong power by large-scale pertaining, how to effectively unleash their power to diverse downstream tasks is still a challenge.
2.3 PEFT in RS
@@
PEFT approaches, such as adapter [28, 27, 31] are proposed to fine-tune large pre-trained models to downstream tasks in Natural Language Processing (NLP) by introducing light-weight learnable parameters. Subsequently, the PEFT has begun to be transferred to vision[69, 1, 39] tasks and rapidly become the focus of researchers. Visual Prompt Tuning (VPT) [37] first introduced ‘prompt tuning’ into the vision area as learnable vectors. The thought of this paradigm is also similar to the adversarial re-programming [15]. Currently, cross-domain works also focus on PEFT, such as [18] which is the first prompt tuning method in Domain Adaptation (DA) [59, 17, 20] and [67] is the first work leveraging PEFT to fine-tuning VFMs for Domain Generalization (DG) [2]. In the RS field, the exploration of PEFT is also vigorous. For example, UPetu [13] addresses storage for dense prediction tasks. TEA [32] uses an adapter network and top-down guidance for detection and segmentation tasks. [53] also leverage LoRA to solve oriented detection task. However, for the cross-domain semantic segmentation tasks, the research of PEFT is relatively blank, and existing methods can not handle the influence of artifacts. Earth-Adapter is the first PEFT for this area and can effectively tackle the artifacts.
3 Method
3.1 Preliminary
@@
We first denote the input RS images as $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$, the corresponding label of $\mathbf{x}$ as $\mathbf{y}\in\mathbb{R}^{H\times W\times K}$, and the semantic segmentation prediction of $\mathbf{x}$ as $\mathbf{y}^{\prime}\in\mathbb{R}^{H\times W\times K}$, where $K$ represents the number of semantic classes. Here, we’ll illustrate our innovative Earth-Adapter overview.
@@
The proposed Earth-Adapter is composed of two key components: the DFT and MoA. MoA consists of a Spatial Adapter, a LF Adapter, a HF Adapter, and a router. We employ $R_{\xi}$, parameterized by $\xi$, to denote the router and $E^{i}_{\epsilon},\;i\in\{1,2,3\}$, parameterized by $\epsilon$, to represent the three adapter experts. For the network architecture, we adopt the framework introduced by Rein [67], integrating DINOv2 [52] as the backbone, represented as $f_{\phi}$ and parameterized by $\phi$, alongside Mask2Former [10] as the decoder, denoted by $f_{\theta}$ and parameterized by $\theta$. Section 3.2 illustrates the Earth-Adapter’s design, while Section 3.3 elaborates on our cross-domain training framework.
3.2 Details of Earth-Adapter
The core idea of Earth-Adapter is the seamless integration of highly efficient, learnable layers with minimal parameters into the frozen FMs.
@@ 3.2 Details of Earth-Adapter
Optimization objective. Before employing the Earth-Adapter, the optimization objective of fine-tuning FMs is to identify a set of parameters that minimize the loss of the entire model on the downstream task:
@@
where $\mathcal{D}$ is the dataset, $f_{\phi}\cdot f_{\theta}$ is the segmentation model, and $\mathcal{L}_{seg}$ is the loss function, here we use the default CE loss. After activating our Earth-Adapter, the optimization objective becomes:
@@
During training, the original parameters of the backbone are kept frozen (denoted as $f_{\phi}^{*}$), ensuring they remain unaltered. In the following description, we denote the segmentation network combined with Earth-Adapter as $G_{\theta,\xi,\epsilon}$.
Parallel Frequency Adaptation. Let $\mathbf{F}_{i}\in\mathbb{R}^{n\times c}$ denote the visual feature at the $i$-th layer of the backbone, where $n$ represents the token sequence length and $c$ denotes the channel dimension. Our frequency adaptation operates through parallel processing streams: The first Spatial Adapter employs a low-rank projection to refine the spatial feature:
@@
where $\Delta\mathbf{F}^{spatial}_{i}$ means the processed i-th layer Spatial-Domain features, and $\textbf{Adapter}_{1}$ represents a nonlinear mapping layer composed of two low-rank matrices:
@@
The Frequency Adapter decompose features through spectral analysis. We first reshape $\mathbf{F}^{spatial}$ to $({C,H,W})$ and apply DFT on spatial features as $\mathcal{FT}(\mathbf{F}_{spatial})$. When splitting the frequency domains, we employ a fixed frequency cutoff $\rho$ to decompose the spectrum into high and low frequency components. Subsequently, these components are transformed back into features via the Inverse Fourier Transform (IFT), yielding LF and HF features:
@@
Mixture of Adapters. We implement dynamic feature fusion through a Router that learns optimal combinations of feature adjustment according to the original visual features. The Router weight is computed by channel-wise attention:
@@
$\Delta\mathbf{F}_{i}=\alpha_{i}\sum_{k=1}^{3}\mathbf{w}^{(k)}\Delta\mathbf{F}^{(k)}_{i},$ (11)
where $\alpha_{i}$ is a learnable scaling parameter with small initial value, and $k\in\{spatial,low,high\}$. The complete feature optimization also maintains residual structure:
@@
3.3 Training FrameWork
In this study, we explore the application of Earth-Adapter across two distinct tasks: DG and UDA Semantic Segmentation. For DG Semantic Segmentation tasks, we train Earth-Adapter with Mask2former in end-to-end supervised learning as Equation 2. For UDA Semantic Segmentation, we use the self-training framework introduced in DACS [61] without using target domain labels. The pseudo labels for target domain images are produced by a teacher network $G_{\theta,\xi,\epsilon}^{\dagger}$:
where []delimited-[][$\cdot$][ ⋅ ] denotes the Iverson bracket, and $j\in HW$ is the spatial index of the pixel in the image. The teacher network remains inactive during the training process, with its parameters being updated based on an exponential moving average (EMA) of the student model’s parameters:
where $(\mathbf{x_{S}},\mathbf{y_{S}})$ is the data sample from the source domain, and $\mathcal{M}$ is generated by randomly selecting half of the categories based on the labels of the source domain $\mathbf{y_{S}}$. The Optimization objective for the UDA task if defined by:
@@ 3.3 Training FrameWork
$\lambda_{uda}$ is the weight parameter for UDA loss.
4 Experiments
Backbone Methods Cross-Domain Adaptation Avg. Cross-Domain Generalization Avg.
@@ 4 Experiments
MiT-B5 [68] DAFormer [29] 64.4 54.8 52.7 42.5 53.6 42.4 41.4 54.2 39.9 44.5
HRDA [30] 67.6 58.6 53.2 35.3 53.7 33.1 31.1 54.2 39.8 39.6
DINOv2-L [52] Frozen 21.0 7.2 21.5 11.8 15.4 57.9 49.4 57.1 42.7 51.8
@@
Adapter [28] 66.4 59.3 55.9 46.2 57.0 56.0 47.4 57.6 41.8 50.7
LoRA [31] 17.8 18.8 24.5 26.0 15.7 20.3 25.6 29.3 21.9 24.3
VPT [37] 66.2 59.3 55.3 48.0 57.2 59.2 52.3 57.4 44.9 53.5
AdaptFormer [7] 12.9 15.3 25.1 22.2 18.9 14.0 19.3 31.0 15.8 20.0
Rein (Baseline) [67] 60.2 60.9 52.8 26.0 50.0 60.8 52.5 55.8 43.4 53.1
Earth-Adapter (Ours) 67.7 (+7.5) 62.2 (+1.3) 55.9 (+3.1) 50.0 (+24.0) 59.0 (+9.0) 64.9 (+4.1) 55.1 (+2.6) 59.0 (+3.2) 45.7 (+2.3) 56.2 (+3.1)
Improvements over the second best +0.1 +1.3 +0.0 +2.0 +1.8 +4.1 +2.5 +1.4 +0.8 +2.7
Table 1: Performance Comparison between representative DA models, PEFT methods, and our Earth-Adapter on DA and DG benchmarks. The Bold font in the table means the SOTA performance. The ‘+’ in the bracket means the improvement compared with our baseline Rein. The last line of data means the improvement compared with the second-best performance of all other methods.
Methods Setting Surf Bldg Vegt Tree Car Clut mIoU (%) Setting Bkgd Bldg Road Water Barr Frst Agri mIoU (%)
Frozen P2V (DA) 38.6 24.1 11.9 47.2 2.5 1.5 21.0 U2R (DA) 42.6 17.2 6.7 9.6 0.7 5.7 0.0 11.8
Rein [67] 75.4 88.6 50.3 73.4 59.0 14.5 60.2 27.9 41.9 29.3 51.2 0.8 53.2 0.0 29.2
Earth-Adapter (Ours) 80.3 89.5 66.3 74.4 61.9 33.6 67.7 56.3 63.4 48.7 66.6 13.7 38.4 63.2 50.0
Frozen P2V (DG) 70.6 79.9 46.1 72.1 58.6 20.4 57.9 U2R (DG) 56.7 54.8 41.3 67.0 14.6 15.1 49.2 42.7
Rein [67] 74.4 85.1 51.0 23.0 57.1 24.5 60.8 53.0 57.8 41.2 61.1 9.4 23.4 57.8 43.4
Earth-Adapter (Ours) 76.3 85.4 59.0 74.1 54.2 40.6 64.9 52.6 55.7 44.1 64.6 17.0 29.6 56.1 46.7
Table 2: Performance Comparison of Per-Class between Frozen DINOv2-L, Rein, and Earth-Adapter on Potsdam RGB-to-Vaihingen. Both Rein and ours are based on the Frozen DINOv2-L. The Bold font means the best performance.

### v4

# with a Frequency-Guided Mixture of Adapters

@@ Abstract
Vision Foundation Models (VFMs), while powerful, often struggle in Remote Sensing (RS) segmentation tasks when combined with existing Parameter-Efficient Fine-Tuning (PEFT) methods. We observe that this limitation primarily arises from their inability to effectively handle the pervasive artifacts in RS imagery. To address this, we introduce Earth-Adapter, the first PEFT method specifically designed for RS artifact mitigation. Earth-Adapter introduces a novel Frequency-Guided Mixture of Adapters (MoA) approach, structured around a “divide and conquer” strategy. It first utilizes Discrete Fourier Transformation (DFT) to ”divide” features into distinct frequency components, thereby effectively isolating artifact-related information from semantic signals. Subsequently, to “conquer” these artifacts, MoA independently optimizes features within different subspaces and dynamically assigns weights via a router to aggregate the refined representations. This enables adaptive refinement of the VFM’s representation space to mitigate the impact of artifacts. This simple yet highly effective PEFT method demonstrably mitigates artifacts and significantly enhances VFMs’ performance on RS segmentation tasks. Extensive experiments demonstrate Earth-Adapter’s effectiveness on in-domain semantic segmentation (SS), as well as Domain Adaptive (DA) and Domain Generalized (DG) semantic segmentation tasks. Compared with the baseline Rein, Earth-Adapter significantly improves mIoU by 1.2% in SS, 9.0% in DA, and 3.1% in DG benchmarks. Our code is available at https://github.com/VisionXLab/Earth-Adapter.
Introduction
Figure 1: Performance across various remote sensing image segmentation benchmarks between Frozen VFM (DINOv2-L), Rein (Baseline) and the proposed Earth-Adapter.
Vision Foundation Models (VFMs) have made significant strides in recent years, driven by both natural language supervision paradigms (e.g., CLIP (Radford et al. 2021), ALIGN (Jia et al. 2021), SigLIP (Zhai et al. 2023)) and self-supervised paradigms (e.g., MAE (He et al. 2022a), DINOv2 (Oquab et al. 2023)). Pre-training on ultra-large-scale datasets empowers VFMs with robust zero-shot comprehension capabilities, leading to their extensive adoption across diverse tasks. These include classical computer vision tasks (classification (Zhou et al. 2022), detection (Zareian et al. 2021; Minderer et al. 2022), segmentation (Chen et al. 2024a)) as well as multimodal learning (Liu et al. 2023). Simultaneously, the conventional pre-train-then-fine-tune paradigm is evolving. When adapting VFMs to downstream tasks, the key challenge becomes how to efficiently preserve and unleash their inherent capabilities. Under this context, Parameter-Efficient Fine-Tuning (PEFT) methods (Hu et al. 2021; Jia et al. 2022; Houlsby et al. 2019a; Gong et al. 2024) emerge as the pivotal solution due to their superior parameter-performance trade-off.
@@ Introduction
Existing PEFT methods designed for language or nature imagery have made great progress with excellent works, such as LoRA (Hu et al. 2021), Visual Prompt Tuning (VPT) (Jia et al. 2022). However, most exhibit significant performance degradation in remote sensing (RS) segmentation tasks when integrated with DINOv2 (Oquab et al. 2023) (experimental evidence provided in later sections). We argue the main reason lies in the influences of artifacts in the features. As shown in Figure 2 (a), we show the PCA visualization of features of DINOv2-L, which contains obvious redundant artifacts. We have also observed that artifacts in natural images exhibit distinct characteristics compared to those in RS images. In natural images, artifacts typically surround foreground objects (Darcet et al. 2024), such as humans or animals, and the disturbances they cause are relatively limited. In contrast, RS images, due to their overhead perspective, lack centralized subjects and contain multiple coexisting multi-scale targets. For example, a single RS image may simultaneously include large-scale agricultural regions and fragmented road networks. As a result, artifacts are almost situated everywhere in RS images shown in Figure 2 (a), causing severe interference to pixel-level feature extraction, which is pivotal for segmentation tasks.
@@
To address artifacts and improve VFM performance in RS segmentation tasks, we propose a “divide-and-conquer” strategy. In the “divide” stage, based on the observation that high-frequency (HF) signals capture local details while low-frequency (LF) signals encode global structures, we apply Discrete Fourier Transformation (DFT) to separate HF and LF components, isolating artifacts in the HF domain. In the “conquer” stage, we design the Mixture of Adapters (MoA), which adjusts features in their respective frequency domains and then uses a router to dynamically assign weights for aggregating the corrected features. These aggregated features are further combined with the original frozen VFM features through skip connections and dynamic scaling coefficients, preserving the strong feature extraction ability of the underlying model.
@@
This unique pipeline forms Earth-Adapter, a PEFT method tailored for RS semantic segmentation. We comprehensively evaluate Earth-Adapter on 12 established benchmarks across three settings—in-domain (SS), domain adaptive (DA), and domain generalized (DG)—and demonstrate its effectiveness. Compared with existing PEFT methods, Earth-Adapter achieves state-of-the-art (SOTA) performance on SS, DA and DG benchmarks and outperforming our baseline Rein (Wei et al. 2024) by 1.2% , 9.0% and 3.1% mIoU, respectively. We also conduct an in-depth analysis of Earth-Adapter’s design and potential. The core contributions can be summarized as follows:
  • We observe that VFMs are prone to artifact interference in the high-dimensional feature space of RS imagery, which hinders their adaptation to RS segmentation tasks.
  • To address this, we propose Earth-Adapter, the first PEFT approach tailored specifically to mitigate artifact-related issues.
  • We develop a Frequency-Guided Mixture of Adapters (MoA) that effectively isolates artifacts within distinct frequency subspaces, coupled with a dynamic router that adaptively fuses adapter outputs to optimize pixel-level VFM features for RS imagery.
  • Extensive experiments across diverse RS segmentation benchmarks demonstrate the effectiveness of Earth-Adapter.
Related Work
RS Semantic Segmentation
@@
Semantic segmentation (Long, Shelhamer, and Darrell 2015; Shelhamer, Long, and Darrell 2016; Chen et al. 2017) aims to classify every pixel in an image. RS segmentation is a critical task in numerous real-world applications, including land cover mapping, urban planning, and environmental monitoring (He et al. 2022b). Cross-domain task is crucial due to the diversity of data, and lots of DA methods have been applied to improve the generalization ability of segmentation models(Rui and Jintao 2020; Tong et al. 2020). Compared with DA, the number of DG research (Iizuka, Xia, and Yokoya 2023; Liang et al. 2024) in RS is much less than DA. Although these works (Zhu et al. 2023) make contributions, most of them only focus on specialized models rather than leveraging the power of VFMs, leading to limited cross-domain capabilities.
Vision Foundation Models
Vision Foundation Models (VFMs) have made significant progress and have exerted considerable influence across various AI-related fields. Among them, vision-language foundation models—such as CLIP (Radford et al. 2021), ALIGN (Jia et al. 2021), SigLIP (Zhai et al. 2023), and InternViT (Chen et al. 2024b)—are trained on billions of image-text pairs to learn joint visual-textual representations. These models have been widely applied to a broad range of vision and multimodal tasks, including classification (Zhou et al. 2022), detection (Zareian et al. 2021; Minderer et al. 2022), segmentation (Chen et al. 2024a), image-text retrieval (Radford et al. 2021), image captioning (Dai et al. 2023), and large multimodal models (LMMs) (Luo et al. 2024). On the other hand, self-supervised models can also learn rich visual representations from large-scale unlabeled data. MAE (He et al. 2022a) learns generalized visual features by reconstructing masked images from partial inputs. MoCo (Chen, Xie, and He 2021) employs a momentum encoder and contrastive learning to extract discriminative representations. DINOv2 (Oquab et al. 2023) combines self-distillation with contrastive learning to produce scalable, transferable features for various downstream tasks. Beyond these various VFMs, a key challenge remains in effectively unlocking their potential to enhance performance on downstream tasks. Earth-Adapter is a pioneering work for RS segmentation tasks
PEFT in RS
@@
PEFT approaches, such as adapter (Houlsby et al. 2019a, b; Hu et al. 2021) are proposed to fine-tune large pre-trained models to downstream tasks in Natural Language Processing (NLP) by introducing light-weight learnable parameters. Subsequently, the PEFT has begun to be transferred to vision(Yin et al. 2023; Agiza, Neseem, and Reda 2024) tasks and rapidly become the focus of researchers. Visual Prompt Tuning (VPT) (Jia et al. 2022) first introduced ‘prompt tuning’ into the vision area as learnable vectors. The thought of this paradigm is also similar to the adversarial re-programming (Elsayed, Goodfellow, and Sohl-Dickstein 2018). Currently, cross-domain works also focus on PEFT, such as (Ge et al. 2023) which is the first prompt tuning method in Domain Adaptation (DA) (Sun et al. 2023; Gao et al. 2022; Gong et al. 2024) and (Wei et al. 2024) is the first work leveraging PEFT to fine-tuning VFMs for Domain Generalization (DG) (Bi et al. 2024). In the RS field, the exploration of PEFT is also vigorous. For example, UPetu (Dong, Gu, and Liu 2024) addresses storage for dense prediction tasks. TEA (Hu et al. 2024) uses an adapter network and top-down guidance for detection and segmentation tasks. (Pu and Xu 2025) also leverage LoRA to solve oriented detection task. However, for the RS semantic segmentation tasks, the research of PEFT is relatively blank, and existing methods can not handle the influence of artifacts. Earth-Adapter is the first PEFT designed to tackle artifacts, thereby unlocking the potential of VFMs for remote sensing segmentation tasks.
Method
Preliminary
@@
We first denote the input RS image as $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$, the corresponding label of $\mathbf{x}$ as $\mathbf{y}\in\mathbb{R}^{H\times W\times K}$, and the semantic segmentation prediction of $\mathbf{x}$ as $\mathbf{y}^{\prime}\in\mathbb{R}^{H\times W\times K}$, where $K$ represents the number of semantic classes.
Here, we’ll illustrate our innovative Earth-Adapter overview.
@@
The proposed Earth-Adapter is composed of two key components: the MoA and Router. MoA consists of a Spatial Adapter, a low-frequency (LF) Adapter, a high-frequency (HF) Adapter. We employ $R_{\xi}$, parameterized by $\xi$, to denote the router and $E^{i}_{\epsilon},\;i\in\{1,2,3\}$, parameterized by $\epsilon$, to represent the three adapter experts. For the network architecture, we adopt the framework introduced by Rein (Wei et al. 2024), integrating DINOv2-L (Oquab et al. 2023) as the backbone, represented as $f_{\phi}$ and parameterized by $\phi$, alongside Mask2Former (Cheng et al. 2022) as the decoder, denoted by $f_{\theta}$ and parameterized by $\theta$. The following presents the details of Earth-Adapter, and we will elaborate on our training framework in the Training Framework section.
Details of Earth-Adapter
Earth-Adapter aims to adapt VFMs to RS image semantic segmentation tasks with minimal learnable parameters.
@@
Optimization objective. Before employing the Earth-Adapter, the optimization objective of fine-tuning VFMs is to identify a set of parameters that minimize the loss of the entire model on the downstream task:
@@
where $\mathcal{D}$ is the dataset and $\mathcal{L}_{seg}$ is the loss function, here we use the default CE loss. After incorporating our Earth-Adapter, the optimization objective becomes:
@@
During training, the original parameters of the backbone (VFM) are kept frozen (denoted as $f_{\phi}^{*}$). In the following description, we denote the segmentation network combined with Earth-Adapter as $G_{\theta,\xi,\epsilon}$.
Mixture of Adapters. Let $\mathbf{F}_{i}\in\mathbb{R}^{(hw)\times c}$ denote the visual feature at the $i$-th layer of the backbone, where $hw$ represents the token sequence length and $c$ denotes the token dimension. Our frequency adaptation operates through parallel processing streams: The first Spatial Adapter employs a low-rank projection to refine the spatial feature:
@@
where $\Delta\mathbf{F}^{spatial}_{i}$ means the processed i-th layer Spatial-Domain features, and $\textbf{Adapter}_{1}$ represents a nonlinear mapping layer composed of two low-rank matrices and an activation:
@@
The Frequency Adapters consist of a HF Adapter (high-frequency subspace) and a LF Adapter (low-frequency subspace), which fine-tune features in specific frequency subspaces derived from a 2D Discrete Fourier Transform (DFT) decomposition. We first reshape $\mathbf{F}^{spatial}$ to (C,H,W)({C,H,W}) and apply DFT on spatial features as $\mathcal{FT}(\mathbf{F}_{spatial})$. When splitting the frequency domains, we employ a fixed frequency cutoff $\rho$ to decompose the spectrum into high and low frequency components. Subsequently, these components are transformed back into features via the Inverse Fourier Transform (IFT), yielding LF and HF features:
@@
The Frequency Adapters shares the same structure as the Spatial Adapter, as defined in Equation 4.
Dynamic Router. We implement dynamic feature aggregation through a Router that learns optimal combinations of feature adjustment according to the original visual features. The Router weight is computed by channel-wise attention:
@@
$\Delta\mathbf{F}_{i}=\alpha_{i}\sum_{k=1}^{3}\mathbf{w}_{i}^{(k)}\Delta\mathbf{F}^{(k)}_{i}$, (11)
where $\alpha_{i}$ is a learnable scaling parameter with small initial value, and $k\in\{spatial,low,high\}$. The frozen features and the refined features are fused via a skip connection:
@@
$\bar{\mathbf{F}_{i}}$ is then forwarded to the subsequent Transformer block to continue the layer-wise processing.
Training Framework
In this work, we explore the application of Earth-Adapter across three distinct semantic segmentation subtasks: semantic segmentation (SS), domain adaptive (DA) semantic segmentation and domain generalized (DG) semantic segmentation . For SS and DGSS, we train Earth-Adapter with Mask2former (Cheng et al. 2022) in end-to-end supervised learning as described in Equation 2. For DA semantic segmentation, we use the self-training framework introduced in DACS (Tranheden et al. 2021) without using target domain labels. DACS employs an EMA teacher to generate pseudo labels for target-domain images, mixes target and source samples to obtain the mixed domain $(\mathbf{x}_{mix},\mathbf{y}_{mix})$ and then jointly trains on both the source and mixed domains. The optimization objective for the DASS if defined by:
@@
$\quad+\lambda_{da}\mathcal{L}_{\text{seg}}\left(G_{\theta,\xi,\epsilon}(\mathbf{x}_{mix}),\mathbf{y}_{mix}\right).$
$\lambda_{da}$ is the weight parameter for DA loss.
Experiments
Methods DASS Avg. DGSS Avg.
@@
Previous SOTA methods
DAFormer 64.4 54.8 52.7 42.5 53.6 42.4 41.4 54.2 39.9 44.5
HRDA 67.6 58.6 53.2 35.3 53.7 33.1 31.1 54.2 39.8 39.6
PEFT-based methods with VFM
Frozen 21.0 7.2 21.5 11.8 15.4 57.9 49.4 57.1 42.7 51.8
Frozen (with register) 11.3 28.2 36.4 16.1 23.0 59.0 51.1 58.9 44.9 53.5
@@
Adapter 66.4 59.3 55.9 46.2 57.0 56.0 47.4 57.6 41.8 50.7
LoRA 17.8 18.8 24.5 26.0 15.7 20.3 25.6 29.3 21.9 24.3
VPT 66.2 59.3 55.3 48.0 57.2 59.2 52.3 57.4 44.9 53.5
AdaptFormer 12.9 15.3 25.1 22.2 18.9 14.0 19.3 31.0 15.8 20.0
Rein (baseline) 60.2 60.9 52.8 26.0 50.0 60.8 52.5 55.8 43.4 53.1
Earth-Adapter (Ours) 67.7 62.2 55.9 50.0 59.0 64.9 55.1 59.0 45.7 56.2
$\Delta$ over baseline method +7.5 +1.3 +3.1 +24.0 +9.0 +4.1 +2.6 +3.2 +2.3 +3.1
Table 1: Performance (mIoU%) comparison between previous SOTA methods and PEFT-based methods on DA and DG benchmarks. Bold indicates the best performance. Underlined results denote the second-best. The last row shows improvements over the baseline.
Method P V L i Avg.
Traditional Methods
DeepLabV3+ (R-101) 74.8 69.7 51.4 53.4 62.3
PSPNet (R-101) 74.7 69.2 48.5 57.5 62.5
UperNet (Swin-B) 75.5 69.4 55.6 67.9 67.1
Segformer (MiT-B5) 75.3 68.4 55.5 66.8 66.5
Mask2Former (Swin-B) 75.7 71.2 54.7 64.0 66.4
PEFT-based methods with VFM
Frozen 75.6 70.4 54.2 67.3 66.9
Frozen (with register) 75.7 70.8 56.0 67.5 67.5
Full Fine-Tune 70.5 64.6 43.9 12.0 47.7
LoRA 71.9 64.6 51.0 48.5 59.0
VPT 75.9 70.6 54.6 68.5 67.4
AdaptFormer 65.4 56.0 42.0 7.5 42.8
Rein 76.2 70.8 54.9 68.4 67.6
Earth-Adapter (Ours) 76.7 71.7 56.9 69.8 68.8
$\Delta$ over baseline method +0.5 +0.9 +2.0 +1.4 +1.2
Table 2: Performance (mIoU%) comparison between previous SOTA methods and PEFT-based methods on $S$ benchmarks. Bold indicates the best performance. Underlined results denote the second-best. The last row shows improvements over the baseline.
General VFMs